OpenStudy (anonymous):

A machine starts dumping sand at the rate of 20 m3/min, forming a pile in the shape of a cone. The height of the pile is always twice the length of the base diameter.After 5 minutes, how fast is the area of the base increasing? When after 5 min the height is 11.516, after 5 min the height is increasing at .7678m/min, after 5 min the base radius is increasing at .19196 m/min

OpenStudy (mathmale):

Hello, S. W.! Nothing like doing a little calculus bright and early on a fine Sunday morning, eh? Have you done this type of problem before? If so, what's the topic? Looking at your problem statement, the following to-do list comes to mind: 1) Write the equations of the volume of a cone and that of the area of the base. 2) Write a formula for the height of the pile in terms of the base radius (not diameter). Base it upon "The height of the pile is always twice the length of the base diameter." 3) Re-write the formula for the volume in terms of the base radius only. 4) Determine the proper label for the quantity described here: "A machine starts dumping sand at the rate of 20 m3/min." 5) Determine the proper label for "how fast is the area of the base increasing" There's more to this process, but my list should get you started.

OpenStudy (anonymous):

Hello, I found that After 5 minutes, how fast is the area of the base increasing? When after 5 min the height is 11.516, after 5 min the height is increasing at .7678m/min, after 5 min the base radius is increasing at .19196 m/min, so I already workd out the volume formula v=1/3pir^2h to solve these

OpenStudy (anonymous):

It is rate of change

OpenStudy (mathmale):

Have you any remaining questions or concerns about this problem? If so, please be specific about what you still need to know to solve it completely.

OpenStudy (anonymous):

Yes I need to figure out how to find "after 5 minutes, how fast is the area of the base increasing"

OpenStudy (mathmale):

If you were able to figure out the first part of this problem, the rest should be fairly straightforward, and should draw upon the work you've already done. At this point I'd ask you to explain what you've done so far and where you're unsure of what to do next. While you don't absolutely have to answer my questions (above), doing so might be useful.

OpenStudy (mathmale):

Re: "It is rate of change:" I'd also call this a "related rates" problem.

OpenStudy (anonymous):

This is what I have done so far 1/ What is the height of a cone volume 100m^3 with a base diameter 1/2 the height? h = 4r V = 1/3 pi r^2*h 100 = 1/3 pi r^2*4r r^3 = 100/(4/3 pi) r = cuberoot(100/(4/3 pi)) 4r = 4*cuberoot(100/(4/3 pi)) h = 11.5176 m 2/ V = 1/3 pi (1/4 h)^2*h = 1/3 pi 1/16 h^3 dV/dt = pi 1/16 h^2 dh/dt 20 = pi 1/16 (11.5176)^2 dh/dt dh/dt = 20/(pi 1/16 (11.5176)^2) = .7678 m/min 3/ As h = 4r Then dr/dt = 1/4 dh/dt dr/dt = .7678/4 = .1920 m/min

OpenStudy (mathmale):

Thank you for sharing this considerable amount of work. Having done this much work, where do you need help?

OpenStudy (mathmale):

I don't see a formula for the area of the base in what you've shared with me. Care to write that formula? Care to differentiate that formula with respect to t?

OpenStudy (anonymous):

Ok so in respect to h if the radius is 1/4*h the area of the base would be a=pi(1/4*h)^2?

OpenStudy (anonymous):

Why in respect to t?

OpenStudy (mathmale):

Let me ask you this: Why did you find the derivatives with respect to t here? " Then dr/dt = 1/4 dh/dt "

OpenStudy (anonymous):

I was looking for a time rate

OpenStudy (anonymous):

soo.. area might be pi 1.4^2 dh/dt

OpenStudy (mathmale):

Finding the derivative of the formula for the area of a circle with respect to t gives you precisely that: the time rate of change of the area with respect to t.

OpenStudy (mathmale):

1. write the formula for the area of the base in terms of the radius, r. 2. differentiate this formula with respect to t. 3. plug in the known values. Identify any values that are not yet known; find them.

OpenStudy (anonymous):

1. it would be pi(1/4h)^2 because radius is in respect h right?

OpenStudy (anonymous):

Well r is cubrt(100/(4/3pi) so I guess the area in respect to r is pi[100/(4/3pi)]^2?

OpenStudy (mathmale):

That looks reasonable. However, the question is not what the area is, but rather how fast the area is changing with respect to time. Agree with that? If A=Pi*r^2, differentiate both sides with respect to t, please.\[\frac{ d }{ dt }(A=\pi*r^2)\]

OpenStudy (anonymous):

Although instead of doing that couldn't I just substitute the rate of change of the height .7678 so the problem would look like da/dt = pi (1/4 * .7678)^2?

OpenStudy (mathmale):

If I understand your math problem correctly, the part we're focusing on now is "After 5 minutes, how fast is the area of the base increasing? " If you see the phrase, "how fast is the area of the base increasing?" you know, or should know, that we need to evaluate a derivative, and that this derivative is (dA/dt). If \(A=\pi*r^2,\) then \(\frac{ dA }{ dt }=2\pi*r\frac{ dr }{ dt}\).

OpenStudy (mathmale):

If you already know r and (dr/dt), all you have to do is to substitute those two quantities into the above equationf or dA/dt.

OpenStudy (mathmale):

After 5 minutes, what is the radius, r? Have you already calculated that? After 5 minutes, what is the rate of change of the radius with respect to t? Have you already calculated that?

OpenStudy (anonymous):

I have "After 5 minutes, what is the rate of change of the radius"

OpenStudy (anonymous):

and I have r = cuberoot(100/(4/3 pi))

OpenStudy (anonymous):

Although if I know that the height is 11.52 and teh radius in respect to h is 1/4h. Than the starting radius would be 2.8

OpenStudy (mathmale):

Are you saying that "at t=5 minutes, r = cuberoot(100/(4/3 pi)) ? If so, good. To find \(\frac{ dA }{ dt },\) you need \(\frac{ dr }{ dt }\) at the specific time t=5 minutes. Do you have that? If not, please find it.

OpenStudy (mathmale):

It's very important that you keep the two different kinds of formulas separate: 1) formulas that relate your variables (except for t) to each other, such as \(A=\pi*r^2\), and 2) formulas that relate rates of change.

OpenStudy (mathmale):

Stop for a moment and ask y ourself, "At this moment, my goal is to find the ( ? ), and to do that, I need this information: ( ? )."

OpenStudy (anonymous):

my goal is to find how fast the area of the base is increasing

OpenStudy (mathmale):

Right, and to do that, "I need ... "

OpenStudy (anonymous):

for that I need to find r and dr/dt

OpenStudy (anonymous):

dr/dt after 5 minutes is .19196

OpenStudy (mathmale):

"and I have to do that for the specific time, t=5 minutes" (assuming that either or both depends upon the time).l

OpenStudy (mathmale):

Your "dr/dt after 5 minutes is .19196" sounds great (although I am not actually checking your numeric results).

OpenStudy (mathmale):

"and my radius, r, after 5 minutes, is ... "

OpenStudy (anonymous):

if the rate of change is .19196 and the original radius is 2.8 than .19196*5=.9598. Then 2.8 +0.9598 = 3.7598?

OpenStudy (mathmale):

The radius of the base of the conical pile after 5 minutes will be greater than the original radius, since the machine continues to dump sand onto the pile. Do you have a formula for the radius as a function of time, t? If you do, write out that formula:\(r(t) = 2.8 + ???\). Then substitute t=5 (minutes). Once again, I'm not checking your work, but rather asking questions that will help you check your results yourself.

OpenStudy (anonymous):

ummm do you meant something like...dV /dt = (4/3) 3r^2 dr/dt

OpenStudy (mathmale):

I think you've already tackled that. If I'm reading your math problem correctly, your focus right now is to find how fast the area of the base is changing. Is (dV/dt) relevant to that? Why or why not?

OpenStudy (anonymous):

hmmm not dV is for the volume I am looking for dr for the radius

OpenStudy (mathmale):

Let's do a quick review here. Our goal is to determine r and dr/dt at t=5 minutes. \[V _{cone}=\frac{ 1 }{ 3 }\pi*r^2*h\] and h=4r (you already have this result).

OpenStudy (anonymous):

Are you saying do 2.8 +(1/4) dh/dt because dr/dt =(1/4) dh/dt

OpenStudy (mathmale):

Thus, \[V _{cone}=\frac{ 4 }{ 3 }\pi*r^3\]

OpenStudy (mathmale):

Are you completely comfortable with this discussion, so far? If not, where do you need clarification?

OpenStudy (anonymous):

well I was until we got to solving for r

OpenStudy (anonymous):

It didn't seem like another derivative was needed to solve

OpenStudy (mathmale):

OK: So, we have a formula for V in terms of r, and you want to solve for r? \[V=\frac{ 4 }{ 3 }\pi*r^3\rightarrow 3V=4\pi*r^3, OK?\]

OpenStudy (mathmale):

(All I did was to multiply b oth sides of this equation by 3.)

OpenStudy (mathmale):

I'm waiting for y our "OK."

OpenStudy (anonymous):

that isn't finished is it? you must go on to solve for r in that equation?

OpenStudy (mathmale):

Right, you must.

OpenStudy (anonymous):

Ok

OpenStudy (mathmale):

If 3V=4*Pi*r^3, divide both sides of this equation by (4*Pi). What do you get?

OpenStudy (anonymous):

3v/4pi= r^2

OpenStudy (mathmale):

That's fine, except that it'd be r^3, not r^2, right?

OpenStudy (anonymous):

right so it would be cubrt(3c/4pi)= cubrt(r^3)

OpenStudy (mathmale):

You're on the right track. I'd hope that you'd label this quantity:\[r=\sqrt[3]{\frac{ 3V }{ 4\pi}}\]

OpenStudy (mathmale):

Here you have the formula for r that you wanted. To calculate r after 5 minutes, you'll need to know the volume of the cone. Simply substitute that volume into the formula for r, above, and you'll then have the radius of the base of the cone.

OpenStudy (mathmale):

Your final goal is to determine how fast the area of the base is changing after 5 minutes. Since\[A=\pi*r^2,\]

OpenStudy (anonymous):

Wait but for after five minutes (20m^3/min*5= 100m^3/min) I had r = cuberoot(100/(4/3 pi))

OpenStudy (mathmale):

\[\frac{ dA }{dt }=2\pi*r \frac{ dr }{ dt }\]

OpenStudy (mathmale):

You already have a formula for r (see above). I believe we also found a formula for (dr/dt). If so, we're all set. Just substitute your r and your (dr/dt) into the above formula for (dA/dt) and you're done (assuming that you've calculated r and (dr/dt) for time t=5 min.).

OpenStudy (jack1):

\[\large h = 4r\] \[\large V = \pi r^2 \frac h3\] so \[\large V = \pi r^2 \frac {4r}3\] \[\large V = \pi r^3 \frac 43\] when t = 5mins: \[\large V = \frac {20m^3}{min} \times 5mins\] \[\large V = 100m^3\] therefore r = ? \[\large V = \pi r^3 \frac 43\] \[\large 100 = \pi r^3 \frac 43\] \[\large r^3 = \frac {300}{4 \pi}\] \[\large r = \sqrt[3]{75/\pi}\] \[\large h = 4r\] \[\large h = 4 \sqrt[3]{75/\pi}\] base area change as a function of volume change volume is increasing at 20m^3/min if \[\large V = \pi r^3 \frac 43\] then \[\Large r = \sqrt[3]{ \frac {3V}{4 \pi}}\] if volume at 5mins = 100, and volume increases at 20m^3/min \[\large V = 20t\] therefore \[\Large r = \sqrt[3]{ \frac {3\times 20t}{4 \pi}}\] \[\Large r = \sqrt[3]{ \frac {60t}{4 \pi}}\] \[\Large r = \sqrt[3]{ \frac {15t}{\pi}}\] \[\Large \frac{dr}{dt} = r' = \frac {(\frac 5\pi)^{1/3}}{3^{2/3} \times \sqrt[3]{t^2}}? \] i think...?

OpenStudy (mathmale):

@Jack1: Thank you for contributing. I'll leave it to you to compare what you have done, on one side, and what S.W. and I, on the other side, have done. if you find discrepancies, we could discuss and resolve them. Note that S.W. needs to find the actual (numeric) rate of change of the area of the base with respect to time after 5 minutes.

OpenStudy (anonymous):

xx confusald

OpenStudy (mathmale):

@S.W.: I'd suggest you read over our entire conversation and determine whether there's anything left that you're not sure about. I'll give you some time to do that. If you feel confused, it's up to you to explain where and why you feel confused, so that Jack1 or I could address that.

OpenStudy (anonymous):

of course

OpenStudy (mathmale):

I may not answer right away, if and when you reply again, but eventually I'll see your work and respond to your questions/concerns. In the meantime, you might want to move on to another math problem.

OpenStudy (anonymous):

This is the only one

OpenStudy (jack1):

k @S.W. so ur equations and mine match up for the volume and height at t = 5 mins V = 100, h = 11.518, and r = 2.879

OpenStudy (anonymous):

I feel like I had already had r and dr/dt... actually I did... derp. r=2.8 after 5 min and dr/dt= .19196 after five min

OpenStudy (anonymous):

Now it seems as though I just need to substitute. 2pi * 2.9* 0.19196= 3.38

OpenStudy (mathmale):

Great to see you two working together. By continuing to compare your results and to resolve any differences, you'll most likely come up with the correct solution. Good going!

OpenStudy (anonymous):

it seems like dA/dt= 2pir dr/dt would be the circumference

OpenStudy (jack1):

V always = 20 t, where t is the time in minutes therefore \[\Large r = \sqrt[3]{ \frac {15t}{\pi}}\] so 15/pi approximately = 4.7746, therefore \[\Large r = \sqrt[3]{ 4.7746t}\] \[\Large r = 1.6839 t^{1/3}\] \[\large dr/dt = \frac {0.5613}{t^{2/3}}\] rate of change for r = {0.5613} / {t^{2/3} Area of base = pi x r^2 \[Area= pi \times r^2\] \[\large Area= pi \times (1.6839 t^{1/3})^2\] \[\large Area= pi \times 2.83552 t^{2/3}\] therefore base area as a function of time = \[\large Area= 8.908 t^{2/3}\] therefore if r is doubled, area is quadrupled, as area is a function of r^2, not r so area of base is increasing at

OpenStudy (anonymous):

Ok, but I don't need to know what the area of the base is I need to know how fast it is increasing

OpenStudy (jack1):

\[\Large dA/dt (8.908 t^(2/3)) = \frac {5.93867}{t^{1/3}}\]

OpenStudy (jack1):

so that's how fast it's increasing

OpenStudy (anonymous):

ummm some how this should be in m/min

OpenStudy (jack1):

units of this are meters / min, simply plug in your required value for time if u want when t = 5 mins 5.93867/t^1/3 5.93867/5^1/3 =...?

OpenStudy (anonymous):

Hmmm oh I see. Are you sure this is correct?

OpenStudy (jack1):

not sure, checkin now, one sec

OpenStudy (jack1):

my bad, the units are in m^2/minute, not meters/minute otherwise...yep, i'm sure, above is ROC of cones' base area, so when t = 5, instantaneous ROC is 3.4729 m^2/min

OpenStudy (anonymous):

how did you get 3.4729 m^2/min?

OpenStudy (jack1):

dA/dt = 5.93867/t^1/3... so if t = 5 = 5.93867/5^1/3 = 3.4729 m^2/min

OpenStudy (anonymous):

Oh ok I calculated wrong . This answer is very similar that that of dA/dt=2 pi* 2.8*.19192 do you think they correlate?

OpenStudy (jack1):

unsure, where/how did u derive that formula from?

OpenStudy (anonymous):

Actually they are the same answer. @mathmale gave it to me

OpenStudy (jack1):

cools... but importantly do u understand how we got here?

OpenStudy (anonymous):

hahah yes I've been there for the past 3 hours, but my way was not so complex. I am sure your ways were the true mathematical ways to solve

OpenStudy (jack1):

all good dude, so long as u can arrive at the right answer in all situations use ur way if it's easier

OpenStudy (anonymous):

yeah right that wouldn't fly over so well with the teacher, he's straight by the book, it's just that eventually I get confused by the details

OpenStudy (anonymous):

Thanks for the help man!

OpenStudy (jack1):

all good dude, u got there hey, we just confirmed u were right, s'all anytime, slaters man!

OpenStudy (anonymous):

A one line solution using Mathematica 9 for the calculations is attached.